The Carubas have a collection of 28 video movies, including 12 westerns and 16 science fiction. Elise selects 3 of the movies at random to bring to a sleep-over at her friend's house, find the probability of:
P(3 science fiction), P(2 westerns and 1 science fiction) and P(2 science fiction and 2 westerns).

Answer by stanbon(75887)

The Carubas have a collection of 28 video movies, including 12 westerns and 16 science fiction. Elise selects 3 of the movies at random to bring to a sleep-over at her friend's house, find the probability of:
P(3 science fiction) = 16C3/28C3 = 0.1709..
-------------------------
P(2 westerns and 1 science fiction) = [12C2*16C1]/[28C3] = 0.3223...
-----------------------
P(2 science fiction and 2 westerns) = [16C2*12C2]/[28C4] = 0.3868...
Comment: This last selection involves selecting 4 movies.
===================
Cheers,
Stan H.
---------------------------